What is the percent composition of carbon by mass in creatine, C4H9N3O2? 




1
Expert's answer
2020-09-21T06:50:34-0400

M (C4H9N3O2) = 131.13 g/mol

%C = (4Ar(C) / M(C4H9N3O2)) * 100% = (48.04 / 131.13) * 100% = 36.64 %

Answer: 36.64 %